OpenStudy (anonymous):

If f(x) = 5x – 3, then what is f^–1(x) ?

OpenStudy (anonymous):

you mean inverse of f(x)?

OpenStudy (anonymous):



OpenStudy (anonymous):

yup its inverse :D

OpenStudy (anonymous):

sfx x = (1/5)*y+3/5 swap variables f^-1(x)=y= (1/5)*x+3/5

OpenStudy (anonymous):